# Deterministic Dynamic Maximal Matching in Sublinear Update Time

### v1

@@ 5.3.4 Analysis of LPM
Proof of Lemma 5.1(1).
[…]
@@ 6 Our Dynamic Maximal Matching Algorithm
The reader should think of $B=n^{\alpha}$ and $\epsilon=1/n^{\beta}$, where $10<\beta<\alpha<$ 1 are some absolute constants.
@@ 6.1 The Framework
Let $G=(V,E)$ be the input graph undergoing edge updates. We maintain a $(B,(1-\epsilon)B)$-EDCS $H:=(V,E(H))$ of $G$ at all times, as per Theorem 2.2. This incurs an update time of $O\left(\frac{n}{B\epsilon}\right)$. The EDCS $H$ is maintained explicitly, and we can make adjacency-list and adjacency-matrix queries to $H$ whenever we want. It takes $O(1)$ time to get the answer to any such query.
@@
Our first task will be to efficiently maintain a matching $M_{\texttt{base}}\subseteq E_{\texttt{core}}$ in the core subgraph, which we will refer to as the base matching. This base matching will satisfy certain desirable properties (see Lemma 6.6). Let $G_{\texttt{adj}}:=(V_{\texttt{adj}},E_{\texttt{adj}})$ denote the subgraph of $G$ induced by the set of nodes that are unmatched under $M_{\texttt{base}}$, that is, $V_{\texttt{adj}}:=V\setminus V(M_{\texttt{base}})$ and $E_{\texttt{adj}}:=\{(u,v)\in E:u,v\in V_{\texttt{adj}}\}$. We will refer to $G_{\texttt{adj}}$ as the adjunct subgraph of $G$ w.r.t. $M_{\texttt{base}}$. Next, we will show how to efficiently maintain a maximal matching $M_{\texttt{adj}}\subseteq E_{\texttt{adj}}$ in $G_{\texttt{adj}}$, which will be called the adjunct matching (see Lemma 6.14). This will immediately imply that $M_{\texttt{final}}:=M_{\texttt{base}}\cup M_{\texttt{adj}}$ is a maximal matching in $G=(V,E)$; and $M_{\texttt{final}}$ will be the output of our dynamic algorithm.
@@
Organization of the rest of Section 6. In Section 6.2, we present a classification of nodes according to their degrees in $H$, which will be used to define the base and the adjunct matchings. In Sections 6.3 and 6.4, we respectively show how to maintain the matchings $M_{\texttt{base}}$ and $M_{\texttt{adj}}$ (see Lemma 6.6 and Lemma 6.14). Finally, we prove Theorem 1.1 in Section 6.5.
@@ 6.2 Classification of Nodes according to EDCS Degrees
For part (1), note that since $(u,v)\in E(H_{\texttt{init}})$, we have $\mathrm{deg}_{H_{\texttt{init}}}(u)+\mathrm{deg}_{H_{\texttt{init}}}(v)\leq B$ (see Definition 2.1). Furthermore, since $u$ is very-high, we have $\mathrm{deg}_{H_{\texttt{init}}}(u)>\left(\frac{1}{2}+\delta\right)B$. Thus, it follows that $\mathrm{deg}_{H_{\texttt{init}}}(v)\leq B-\mathrm{deg}_{H_{\texttt{init}}}(u)<\left(\frac{1}{2}-\delta\right)B$, and so $v$ must be a low node.
For part (2), note that since $(u,v)\in E(H_{\texttt{init}})$, we have $\mathrm{deg}_{H_{\texttt{init}}}(u)+\mathrm{deg}_{H_{\texttt{init}}}(v)\leq B$ (see Definition 2.1). Next, since $u$ is high, we have $\mathrm{deg}_{H_{\texttt{init}}}(u)\geq\left(\frac{1}{2}+\delta-\epsilon\right)B$. Thus, it follows that $\mathrm{deg}_{H_{\texttt{init}}}(v)\leq B-\mathrm{deg}_{H_{\texttt{init}}}(u)\leq\left(\frac{1}{2}-\delta+\epsilon\right)B$. Accordingly, the node $v$ is either low or almost-low.
For part (3), let $e=(u,v)$, where $u$ is low and $v$ is either low or medium. Then $\mathrm{deg}_{H_{\texttt{init}}}(u)<\left(\frac{1}{2}-\delta\right)B$ and $\mathrm{deg}_{H_{\texttt{init}}}(v)<\left(\frac{1}{2}+\delta-\epsilon\right)B$, and hence $\mathrm{deg}_{H_{\texttt{init}}}(u)+\mathrm{deg}_{H_{\texttt{init}}}(v)<(1-\epsilon)B$. As $H_{\texttt{init}}$ is a $(B,(1-\epsilon)B)$-EDCS of $G_{\texttt{init}}$, it follows that $(u,v)\in E\left(H_{\texttt{init}}\right)$ (see Definition 2.1). ∎
@@ 6.3 Maintaining the Base Matching
We will show how to maintain the matching $M_{\texttt{base}}\subseteq E(H_{\texttt{core}})$, satisfying two conditions. Condition 6.4 specifies which nodes are matched under $M_{\texttt{base}}$, whereas Condition 6.5 ensures the “stability” of the base matching as the input graph $G$ undergoes updates within a given phase.
@@
We can maintain a matching $M_{\texttt{base}}\subseteq E(H_{\texttt{core}})$ satisfying Conditions 6.4 and 6.5 by:
@@
The rest of Section 6.3 is organized as follows. In Section 6.3.1, We first show how to compute the matching $M_{\texttt{most}}\subseteq E_{\texttt{init}}$ at the start of a phase. Next, we define the subgraph $H_{\texttt{hilo}}$ in Section 6.3.2, and summarize some of its key properties. In Section 6.3.3 and Section 6.3.4, we respectively show how to initialize the matchings $M_{\texttt{hilo}}$ and $M_{\texttt{base}}$ at the start of a phase, and how to maintain $M_{\texttt{hilo}}$ and $M_{\texttt{base}}$ throughout the updates within a phase. Finally, we put everything together in Section 6.3.5 and prove Lemma 6.6.
@@ 6.3.1 Computing the matching $M_{\texttt{most}}$ at the start of a phase
At the start of a phase, let $H^{\prime}_{\texttt{init}}$ be the subgraph of $H_{\texttt{init}}$ obtained as follows. Initialize $H^{\prime}_{\texttt{init}}\leftarrow H_{\texttt{init}}$. Next, for each very-high node $v\in V_{\texttt{v-hi}}$, arbitrarily keep deleting edges (in $H^{\prime}_{\texttt{init}}$) incident on $v$ until $\mathrm{deg}_{H^{\prime}_{\texttt{init}}}(v)$ becomes equal to $\left(\frac{1}{2}+\delta\right)B$. By part (1) of Proposition 6.1, this procedure does not delete any edge incident on a medium or high node. Thus, when this procedure terminates, we have $\left(\frac{1}{2}-\delta\right)B\leq\mathrm{deg}_{H^{\prime}_{\texttt{init}}}(v)\leq\left(\frac{1}{2}+\delta\right)B$ for every node $v\in V_{\texttt{hi}}\cup V_{\texttt{med}}$, and $\mathrm{deg}_{H^{\prime}_{\texttt{init}}}(v)<\left(\frac{1}{2}-\delta\right)B$ for every node $v\in V_{\texttt{lo}}=V\setminus(V_{\texttt{hi}}\cup V_{\texttt{med}})$. Next, we set:
@@
Since $B=n^{\alpha}$ and $\epsilon=1/n^{\beta}$, for some absolute constants $10<\beta<\alpha<$ 1 (see the comment after (6)), and $\delta=100\epsilon$ (see (6)), we infer that $\Delta\geq\frac{4}{\kappa}$. Also, from the preceding discussion, it follows that $\Delta$ is an upper bound on the maximum degree in $H^{\prime}_{\texttt{init}}$. We next apply the concerned subroutine (deterministic or randomized) from Lemma 3.1, by setting $G:=H^{\prime}_{\texttt{init}}$, and the parameters $\kappa,\Delta$ as in (7). Let $M$ denote the matching returned by the subroutine from Lemma 3.1. At this point, we set $M_{\texttt{most}}\leftarrow M$.
Because of the way we have defined $\kappa,\Delta$ in (7), it follows that $\left(\frac{1}{2}-\delta\right)B\geq(1-\kappa)\Delta$, and hence $V_{\kappa}\supseteq V_{\texttt{hi}}\cup V_{\texttt{med}}$ (see Lemma 3.1 to recall the definition of $V_{\kappa}$). From Lemma 3.1, we now get:
@@
  • Since $H_{\texttt{init}}$ is a $(B,(1-\epsilon)B)$-EDCS of $G_{\texttt{init}}$, it has maximum degree at most $B$ (see Definition 2.1). So, it takes $O(nB)$ time to compute the subgraph $H^{\prime}_{\texttt{init}}$ of $H_{\texttt{init}}$. After that, we can compute $M_{\texttt{most}}$ either in $\tilde{O}(nB/\delta)$ time deterministically, or in $\tilde{O}(nB\delta)$ time with high probability. This leads to an overall deterministic time complexity of $O(nB)+\tilde{O}(nB/\delta)=\tilde{O}(nB/\delta)$, and an overall randomized time complexity of $O(nB)+\tilde{O}(nB\delta)=\tilde{O}(nB)$.
@@ 7                   $H_{\texttt{init}}^{\prime}\leftarrow H_{\texttt{init}}^{\prime}\setminus\{e\}$
[…]
@@ 6.3.2 The subgraph $H_{\texttt{hilo}}$
In words, $E(H_{\texttt{hilo}})$ is the set of all the edges in $H_{\texttt{core}}$ incident on the high nodes, and $V_{\texttt{hilo}}$ is the set of all nodes that are either hi, or low, or almost-low. Since $E(H_{\texttt{core}})\subseteq E(H_{\texttt{init}})$, part-(2) of Proposition 6.1 implies that every edge in $E(H_{\texttt{hilo}})$ has one endpoint in $V_{\texttt{hi}}$ and the other endpoint in $V_{\texttt{lo}}\cup V_{\texttt{a-lo}}$. Thus, $H_{\texttt{hilo}}$ is a bipartite subgraph of $H_{\texttt{core}}$. We will designate the nodes in $V_{\texttt{hi}}$ as being on the “left”, and the nodes in $V_{\texttt{lo}}\cup V_{\texttt{a-lo}}$ as being on the “right”.
We also define two parameters $X,\gamma$ as in (9), which allow us to derive some useful properties of the subgraph $H_{\texttt{hilo}}$. This sets up the stage so that we can apply Lemma 5.1 on $H_{\texttt{hilo}}$.
@@
For part-(2a), note that $V_{\texttt{sf}}\subseteq V_{\texttt{hi}}$ (see Definition 6.2), and for every node $v\in V_{\texttt{sf}}$ we have
@@
In the above derivation, the first step follows from (8), the second step follows from Definition 6.2, and the last step follows from (9). Similarly, Definition 6.2 implies that $\mathrm{deg}_{H_{\texttt{hilo}}}(v)=\mathrm{deg}_{H_{\texttt{core}}}(v)<\left(\frac{1}{2}+\delta-2\epsilon\right)B=X$ for all nodes $v\in V_{\texttt{dmg}}=V_{\texttt{hi}}\setminus V_{\texttt{sf}}$.
For part-(1), consider the situation at the start of a phase. At this point, we have $V_{\texttt{sf}}=V_{\texttt{hi}}$ (see Definition 6.2 and Corollary 6.3). Thus, using the same argument as in the preceding paragraph, we conclude that $\mathrm{deg}_{H_{\texttt{hilo}}}(v)\geq X$ for all nodes $v\in V_{\texttt{hi}}$. Next, recall the way we set $X$ and $\gamma$ in (9). Since $\delta=100\epsilon$ (see (6)), we get:
@@
Part-(2b) follows from Corollary 6.3 and Definition 6.2.
Part-(2c) follows from the definitions of $H_{\texttt{hilo}}$ (see (8)) and $H_{\texttt{core}}$ (see Definition 6.2). ∎

### v2

@@ 5.3.4 Analysis of LPM
Proof of Lemma˜5.1(1).
It is easy to check that the total update time of LPM is dominated by the time to maintain the ES structure. Choosing $q_{\textrm{ep}}=\sqrt{n\gamma}$ results in a total of at most $1+u/q_{\textrm{ep}}$ epochs (Recall that $u$ is the total number of calls to $\textrm{LPM}.\textrm{Delete}$). Applying Lemma 5.20 yields the bound in Lemma 5.1. ∎
5.3.5 Proof of Lemma 5.17
Recall that we refer to vertices in $R_{\textrm{mark}}$ as marked. Define $L_{i}(v)\subseteq L$ to be the set of all vertices $u\in L$ satisfying the following property: the shortest alternating path (w.r.t. $M$) from $u$, which starts with an unmatched edge $e\in E\setminus M$ incident on $u$ and ends at an $M$-free vertex in $R\setminus R_{\textrm{mark}}$, has length exactly $i$. Define $R_{i}(v)$ to be the vertices matched to $L_{i}(v)$ by $M$: so $|L_{i}(v)|=|R_{i}(v)|$. We define $L_{\leq i}=\bigcup_{j\leq i}L_{j}$ and define $L_{<i}$, $L_{>i}$, and $L_{\geq i}$ analogously.
The key to the proof is the following claim.
Claim 5.21.
Say that $|L_{>i}|\geq\frac{2}{\gamma}|R_{\textrm{mark}}|$. Then, $|L_{i+1}|\geq\frac{\gamma}{2}\cdot|L_{>i}|$.
proof of claim:
We know that $|R_{>i}|=|L_{>i}|$. Define $R^{\prime}=R_{>i}\cup R_{\textrm{mark}}$ and note that by assumption of the claim
$|R^{\prime}|\leq|R_{>i}|(1+\gamma/2)=|L_{>i}|(1+\gamma/2).$ (6)
Now, clearly
$L_{i+1}=L_{>i}-L_{\geq i+2}$
We now upper bound $|L_{\geq i+2}|$. It is easy to see that every vertex in $L_{\geq i+2}$ has the property that all of its edges are incident to $R^{\prime}$. Since $G$ has a $\gamma$-degree gap at $X$, counting the edges implies:
$X\cdot|L_{\geq i+2}|\leq(1-\gamma)X\cdot|R^{\prime}|$
Rearranging and appying Equation˜6 implies
$|L_{\geq i+2}|\leq(1-\gamma)|R^{\prime}|\leq(1-\gamma)(1+\gamma/2)|R_{>i}|\leq(1-\gamma/2)|R_{>i}|=(1-\gamma/2)|L_{>i}|$
This implies that
$|L_{i+1}|=|L_{>i}|-|L_{\geq i+2}|\geq\frac{\gamma}{2}\cdot|L_{>i}|$
Proof of Lemma˜5.17 given Claim˜5.21
Let $\delta=\frac{4}{\gamma}\ln(n)$. We will show that $|L_{>\delta}|\leq\frac{2}{\gamma}|R_{\textrm{mark}}|$, which clearly implies Lemma˜5.17. Say for contradiction that $|L_{>\delta}|>\frac{2}{\gamma}|R_{\textrm{mark}}|$. This clearly implies that $|L_{>j}|>\frac{2}{\gamma}|R_{\textrm{mark}}|$ for any $j<\delta$. But this means that for any $j<\delta$ we can apply Claim˜5.21, which in turn implies that for any $j<\delta$ we have
$|L_{>j}|=|L_{>j+1}|+|L_{j+1}|\geq|L_{>j+1}|+\frac{\gamma}{2}\cdot|L_{>j}|\geq(1+\frac{\gamma}{2})|L_{>j+1}|$
The above equation, combined with the fact that for any $x>1$ we have $(1+1/x)^{2x}>e$, implies:
$|L|=|L_{>0}|\geq|L_{>\delta}|\cdot(1+\frac{\gamma}{2})^{\delta}=|L_{>\delta}|\cdot(1+\frac{\gamma}{2})^{\frac{4}{\gamma}\ln(n)}>e^{\ln(n)}=n,$
which contradicts the definition $|L|=n$.
@@ 6 Our Dynamic Maximal Matching Algorithm
The reader should think of $B=n^{\alpha}$ and $\epsilon=1/n^{\beta}$, where $0<\beta<\alpha<1$ are some absolute constants.
@@ 6.1 The Framework
Let $G=(V,E)$ be the input graph undergoing edge updates. We maintain a $(B,(1-\epsilon)B)$-EDCS $H:=(V,E(H))$ of $G$ at all times, as per Theorem˜2.2. This incurs an update time of $O\left(\frac{n}{B\epsilon}\right)$. The EDCS $H$ is maintained explicitly, and we can make adjacency-list and adjacency-matrix queries to $H$ whenever we want. It takes $O(1)$ time to get the answer to any such query.
@@
Our first task will be to efficiently maintain a matching $M_{\texttt{base}}\subseteq E_{\texttt{core}}$ in the core subgraph, which we will refer to as the base matching. This base matching will satisfy certain desirable properties (see Lemma˜6.6). Let $G_{\texttt{adj}}:=(V_{\texttt{adj}},E_{\texttt{adj}})$ denote the subgraph of $G$ induced by the set of nodes that are unmatched under $M_{\texttt{base}}$, that is, $V_{\texttt{adj}}:=V\setminus V(M_{\texttt{base}})$ and $E_{\texttt{adj}}:=\{(u,v)\in E:u,v\in V_{\texttt{adj}}\}$. We will refer to $G_{\texttt{adj}}$ as the adjunct subgraph of $G$ w.r.t. $M_{\texttt{base}}$. Next, we will show how to efficiently maintain a maximal matching $M_{\texttt{adj}}\subseteq E_{\texttt{adj}}$ in $G_{\texttt{adj}}$, which will be called the adjunct matching (see Lemma˜6.14). This will immediately imply that $M_{\texttt{final}}:=M_{\texttt{base}}\cup M_{\texttt{adj}}$ is a maximal matching in $G=(V,E)$; and $M_{\texttt{final}}$ will be the output of our dynamic algorithm.
@@
Organization of the rest of Section˜6. In Section˜6.2, we present a classification of nodes according to their degrees in $H$, which will be used to define the base and the adjunct matchings. In Sections˜6.3 and 6.4, we respectively show how to maintain the matchings $M_{\texttt{base}}$ and $M_{\texttt{adj}}$ (see Lemma˜6.6 and Lemma˜6.14). Finally, we prove Theorem˜1.1 in Section˜6.5.
@@ 6.2 Classification of Nodes according to EDCS Degrees
For part (1), note that since $(u,v)\in E(H_{\texttt{init}})$, we have $\mathrm{deg}_{H_{\texttt{init}}}(u)+\mathrm{deg}_{H_{\texttt{init}}}(v)\leq B$ (see Definition˜2.1). Furthermore, since $u$ is very-high, we have $\mathrm{deg}_{H_{\texttt{init}}}(u)>\left(\frac{1}{2}+\delta\right)B$. Thus, it follows that $\mathrm{deg}_{H_{\texttt{init}}}(v)\leq B-\mathrm{deg}_{H_{\texttt{init}}}(u)<\left(\frac{1}{2}-\delta\right)B$, and so $v$ must be a low node.
For part (2), note that since $(u,v)\in E(H_{\texttt{init}})$, we have $\mathrm{deg}_{H_{\texttt{init}}}(u)+\mathrm{deg}_{H_{\texttt{init}}}(v)\leq B$ (see Definition˜2.1). Next, since $u$ is high, we have $\mathrm{deg}_{H_{\texttt{init}}}(u)\geq\left(\frac{1}{2}+\delta-\epsilon\right)B$. Thus, it follows that $\mathrm{deg}_{H_{\texttt{init}}}(v)\leq B-\mathrm{deg}_{H_{\texttt{init}}}(u)\leq\left(\frac{1}{2}-\delta+\epsilon\right)B$. Accordingly, the node $v$ is either low or almost-low.
For part (3), let $e=(u,v)$, where $u$ is low and $v$ is either low or medium. Then $\mathrm{deg}_{H_{\texttt{init}}}(u)<\left(\frac{1}{2}-\delta\right)B$ and $\mathrm{deg}_{H_{\texttt{init}}}(v)<\left(\frac{1}{2}+\delta-\epsilon\right)B$, and hence $\mathrm{deg}_{H_{\texttt{init}}}(u)+\mathrm{deg}_{H_{\texttt{init}}}(v)<(1-\epsilon)B$. As $H_{\texttt{init}}$ is a $(B,(1-\epsilon)B)$-EDCS of $G_{\texttt{init}}$, it follows that $(u,v)\in E\left(H_{\texttt{init}}\right)$ (see Definition˜2.1). ∎
@@ 6.3 Maintaining the Base Matching
We will show how to maintain the matching $M_{\texttt{base}}\subseteq E(H_{\texttt{core}})$, satisfying two conditions. Condition˜6.4 specifies which nodes are matched under $M_{\texttt{base}}$, whereas Condition˜6.5 ensures the “stability” of the base matching as the input graph $G$ undergoes updates within a given phase.
@@
We can maintain a matching $M_{\texttt{base}}\subseteq E(H_{\texttt{core}})$ satisfying Conditions˜6.4 and 6.5 by:
@@
The rest of Section˜6.3 is organized as follows. In Section˜6.3.1, We first show how to compute the matching $M_{\texttt{most}}\subseteq E_{\texttt{init}}$ at the start of a phase. Next, we define the subgraph $H_{\texttt{hilo}}$ in Section˜6.3.2, and summarize some of its key properties. In Section˜6.3.3 and Section˜6.3.4, we respectively show how to initialize the matchings $M_{\texttt{hilo}}$ and $M_{\texttt{base}}$ at the start of a phase, and how to maintain $M_{\texttt{hilo}}$ and $M_{\texttt{base}}$ throughout the updates within a phase. Finally, we put everything together in Section˜6.3.5 and prove Lemma˜6.6.
@@ 6.3.1 Computing the matching $M_{\texttt{most}}$ at the start of a phase
At the start of a phase, let $H^{\prime}_{\texttt{init}}$ be the subgraph of $H_{\texttt{init}}$ obtained as follows. Initialize $H^{\prime}_{\texttt{init}}\leftarrow H_{\texttt{init}}$. Next, for each very-high node $v\in V_{\texttt{v-hi}}$, arbitrarily keep deleting edges (in $H^{\prime}_{\texttt{init}}$) incident on $v$ until $\mathrm{deg}_{H^{\prime}_{\texttt{init}}}(v)$ becomes equal to $\left(\frac{1}{2}+\delta\right)B$. By part (1) of Proposition˜6.1, this procedure does not delete any edge incident on a medium or high node. Thus, when this procedure terminates, we have $\left(\frac{1}{2}-\delta\right)B\leq\mathrm{deg}_{H^{\prime}_{\texttt{init}}}(v)\leq\left(\frac{1}{2}+\delta\right)B$ for every node $v\in V_{\texttt{hi}}\cup V_{\texttt{med}}$, and $\mathrm{deg}_{H^{\prime}_{\texttt{init}}}(v)<\left(\frac{1}{2}-\delta\right)B$ for every node $v\in V_{\texttt{lo}}=V\setminus(V_{\texttt{hi}}\cup V_{\texttt{med}})$. Next, we set:
@@
Since $B=n^{\alpha}$ and $\epsilon=1/n^{\beta}$, for some absolute constants $0<\beta<\alpha<1$ (see the comment after (7)), and $\delta=100\epsilon$ (see (7)), we infer that $\Delta\geq\frac{4}{\kappa}$. Also, from the preceding discussion, it follows that $\Delta$ is an upper bound on the maximum degree in $H^{\prime}_{\texttt{init}}$. We next apply the concerned subroutine (deterministic or randomized) from Lemma˜3.1, by setting $G:=H^{\prime}_{\texttt{init}}$, and the parameters $\kappa,\Delta$ as in (8). Let $M$ denote the matching returned by the subroutine from Lemma 3.1. At this point, we set $M_{\texttt{most}}\leftarrow M$.
Because of the way we have defined $\kappa,\Delta$ in (8), it follows that $\left(\frac{1}{2}-\delta\right)B\geq(1-\kappa)\Delta$, and hence $V_{\kappa}\supseteq V_{\texttt{hi}}\cup V_{\texttt{med}}$ (see Lemma˜3.1 to recall the definition of $V_{\kappa}$). From Lemma˜3.1, we now get:
@@
  • Since $H_{\texttt{init}}$ is a $(B,(1-\epsilon)B)$-EDCS of $G_{\texttt{init}}$, it has maximum degree at most $B$ (see Definition˜2.1). So, it takes $O(nB)$ time to compute the subgraph $H^{\prime}_{\texttt{init}}$ of $H_{\texttt{init}}$. After that, we can compute $M_{\texttt{most}}$ either in $\tilde{O}(nB/\delta)$ time deterministically, or in $\tilde{O}(nB\delta)$ time with high probability. This leads to an overall deterministic time complexity of $O(nB)+\tilde{O}(nB/\delta)=\tilde{O}(nB/\delta)$, and an overall randomized time complexity of $O(nB)+\tilde{O}(nB\delta)=\tilde{O}(nB)$.
@@ 7       $H_{\texttt{init}}^{\prime}\leftarrow H_{\texttt{init}}^{\prime}\setminus\{e\}$
Algorithm 2 Computing $M_{\texttt{most}}$. We run $\texttt{Initialize-}M_{\texttt{most}}()$ at the start of a phase.
@@ 6.3.2 The subgraph $H_{\texttt{hilo}}$
In words, $E(H_{\texttt{hilo}})$ is the set of all the edges in $H_{\texttt{core}}$ incident on the high nodes, and $V_{\texttt{hilo}}$ is the set of all nodes that are either hi, or low, or almost-low. Since $E(H_{\texttt{core}})\subseteq E(H_{\texttt{init}})$, part-(2) of Proposition˜6.1 implies that every edge in $E(H_{\texttt{hilo}})$ has one endpoint in $V_{\texttt{hi}}$ and the other endpoint in $V_{\texttt{lo}}\cup V_{\texttt{a-lo}}$. Thus, $H_{\texttt{hilo}}$ is a bipartite subgraph of $H_{\texttt{core}}$. We will designate the nodes in $V_{\texttt{hi}}$ as being on the “left”, and the nodes in $V_{\texttt{lo}}\cup V_{\texttt{a-lo}}$ as being on the “right”.
We also define two parameters $X,\gamma$ as in (10), which allow us to derive some useful properties of the subgraph $H_{\texttt{hilo}}$. This sets up the stage so that we can apply Lemma˜5.1 on $H_{\texttt{hilo}}$.
@@
For part-(2a), note that $V_{\texttt{sf}}\subseteq V_{\texttt{hi}}$ (see Definition˜6.2), and for every node $v\in V_{\texttt{sf}}$ we have
@@
In the above derivation, the first step follows from (9), the second step follows from Definition˜6.2, and the last step follows from (10). Similarly, Definition˜6.2 implies that $\mathrm{deg}_{H_{\texttt{hilo}}}(v)=\mathrm{deg}_{H_{\texttt{core}}}(v)<\left(\frac{1}{2}+\delta-2\epsilon\right)B=X$ for all nodes $v\in V_{\texttt{dmg}}=V_{\texttt{hi}}\setminus V_{\texttt{sf}}$.
For part-(1), consider the situation at the start of a phase. At this point, we have $V_{\texttt{sf}}=V_{\texttt{hi}}$ (see Definition˜6.2 and Corollary˜6.3). Thus, using the same argument as in the preceding paragraph, we conclude that $\mathrm{deg}_{H_{\texttt{hilo}}}(v)\geq X$ for all nodes $v\in V_{\texttt{hi}}$. Next, recall the way we set $X$ and $\gamma$ in (10). Since $\delta=100\epsilon$ (see (7)), we get:
@@
Part-(2b) follows from Corollary˜6.3 and Definition˜6.2.
Part-(2c) follows from the definitions of $H_{\texttt{hilo}}$ (see (9)) and $H_{\texttt{core}}$ (see Definition˜6.2). ∎